Verified Answer :
Since there is replacement, drawings are independent which means that the result of one of them does not affect the others.
Probability of drawing an apple is 4/10 (4 apple out of 10 fruits)
Probability of drawing an apricot is 2/10
Probability on one happening consecutively to the other (thanks to independence) is 4/10*2/10=8/100
Which can be simplified to 2/25
Probability of getting an apple then apricot is option (B) \(\frac{2}{25}\)
What is probability?
Probability is the branch of mathematics concerning numerical descriptions of how likely an event is to occur.
Number of apple in the bag = 4
Number of plum in the bag = 1
Number of apricots in the bag = 2
Number of orange in the bag = 3
Total number of fruits in the bag = 10
Here a piece of fruits are drawn twice with replacement
Therefore first event and second events are independent
Probability = Number of favorable outcome / Total outcome
Probability of getting apple = \(\frac{4}{10} =\frac{2}{5}\)
Probability of getting apricot = \(\frac{2}{10} =\frac{1}{5}\)
Probability of getting apple then apricot = Probability of getting apple × Probability of getting apricot
Probability of getting apple then apricot= \((\frac{2}{5})(\frac{1}{5})=\frac{2}{25}\)
Hence, the probability of getting an apple then apricot is option (B) \(\frac{2}{25}\)
Learn more about probability here
https://brainly.com/question/11234923
#SPJ2